Solve the following initial value problem. St/2 if 0 <t<6 y +y= 3 ift > 6 6 y(0) = y(0) = 0 14Pm1011* 1917 Prid A++ V Top1
Answer #1

t12 y+y = { ost/6 it t26 y (0) = 4 (0) = 0 3 fo ost26 y+4 = 1/2 mt tor y+y=0 Let y=e mt y sme y sme smamt SO, mt mt te meSO + 6 yft = C, cost + s sint + 3 (t is conti et diff So, YC6-) - 2 sin 6 +3=4 (6+) = C, Cus6+q simb +3 yC6-) - -), C056+sC, cos 6 = 1/₂ coso sing C = -1/₂ sing oct 26 So, - sirt +ty, tz6 -\y sin 6 cost + (-s + ky cos 6) sint +3 th sint + 3 - sin